Using a FanIn/Out module from near BigBite, an ADC gate width problem was corrected.
At the same time, a 50ns ADC gate was created for BPM. That gate uses cable 07, which feeds into the front of the 1881 card in slot 25 of ROC2.
Further, the L1A coming in from the BigBite weldment was split (using that same module), with one output going to the RT module and one going to the scalers/TDCs.
The scaler map for the right side was updated (with the copy in $DB_DIR to be updated soon, after corrections to the left side are made as well).
The 100kHz signal from the left arm is missing, but that should be corrected shortly.
A followup showing the TDC map for the 1877 with the triggers will be posted soon.